# Geodesics in the Deep Linear Network

### v1

@@ Abstract.
We derive a general system of ODEs and associated explicit solutions in a special case for geodesics between full rank matrices in the deep linear network geometry. In the process, we characterize all horizontal straight lines in the invariant balanced manifold that remain geodesics under Riemannian submersion.
@@ 1. Introduction
The deep linear network (DLN) is a phenomenological model of the training dynamics of neural networks [1, 2, 6, 9]. Given a depth $N$ and width $d$, in its simplest form the DLN consists of the “upstairs” (high-dimensional) parameter space of NN $d\times d$ matrices $\mathbb{M}_{d}^{N}$ and a “downstairs” (low-dimensional) observable space $\mathbb{M}_{d}$, linked by the product map $\phi:\mathbb{M}_{d}^{N}\to\mathbb{M}_{d}$ that sends
@@ 1.2. Balancedness and Riemannian Geometry
Interestingly, the DLN gradient flow (Equation 1.2) has invariant manifolds with associated metrics that illustrate the geometric structure underlying the DLN. One such invariant manifold is defined by the constraints
@@
for all $1\leq p\leq N-1$ [2, Theorem 1]. We further restrict to when all coordinates are rank $d$ and denote this manifold $\mathcal{M}$, the balanced manifold. When $\mathbf{W}\in\mathcal{M}$, $X=\phi(\mathbf{W})\in\mathfrak{M}_{d}$.
It is not hard to show that in $\mathcal{M}$, $W_{p}$ have the same singular values for all $1\leq p\leq N$ and the right singular vectors of $W_{p+1}$ align with the left singular vectors of $W_{p}$ for $1\leq p\leq N-1$. Thus, given the SVD of a matrix $X=Q_{N}\Sigma Q_{0}$, we can parameterize any $\mathbf{W}\in\mathcal{M}$ such that $\phi(\mathbf{W})=X$ using just $(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})$, where $\Lambda=\Sigma^{\frac{1}{N}}$ and $W_{p}=Q_{p}\Lambda Q_{p-1}^{T}$ are obtained by the SVD for all $1\leq p\leq N$. Therefore, there exists a bijective map
@@
$g^{N}$ is crucial in the DLN because it can be used to express the training dynamics of the end-to-end matrix $X$ as a Riemannian gradient flow [3, Section 4]. Furthermore, these metrics in tandem provide a geometric perspective on the DLN model from [10, Theorem 8] and [11, Theorem 36], which we elaborate on in Section 3.4.1:
@@
[5, Section 4] derive explicit formulas for geodesics in the Bures-Wasserstein geometry by identifying a Riemannian submersion structure from invertible matrices to the positive definite matrices. We present an alternative proof of the same geodesic formulas that also generalizes to the DLN, yielding the general ODEs for geodesic in $(\mathfrak{M}_{d},g^{N})$. Furthermore, Theorem 1.1 implies that the DLN also has a Riemannian submersion structure with $(\mathcal{M},\iota)$, $(\mathfrak{M}_{d},g^{N})$, and $\phi$, which we leverage to solve the general equations in a specific case.
@@ 2. Main Results
Our main results are a statement of ODEs for general geodesics in $(\mathfrak{M}_{d},g^{N})$ and explicit solutions when the singular vectors of the endpoints are related by a single matrix $Q\in O_{d}$. The solutions are obtained by characterizing horizontal straight lines in $\mathcal{M}$ in a manner similar to the classical theory of confocal quadrics [7].
@@
We discuss the origin of this matrix in Section $3.$
@@
$X(t)=U(t)\Sigma(t)V(t)^{T}$ is a geodesic between $A$ and $B$ on $(\mathfrak{M}_{d},g^{N})$ if it satisfies the Dirichlet boundary value problem
@@
We discover these equations via direct computation in Hamiltonian coordinates (Section 3.3). Using Riemannian submersion, we are then able to derive explicit solutions to these equations for particular endpoints. As an important intermediate result, we characterize the straight lines in $\mathcal{M}$.
Lemma 2.2 (Characterization of Straight Lines on $\mathcal{M}$).
Let $\mathbf{A}=\xi(\Lambda,Q_{N},\ldots,Q_{0})$ and $\mathbf{B}=\xi(\tilde{\Lambda},\tilde{Q}_{N},\ldots,\tilde{Q}_{0})$ be the endpoints of interest. Let
@@
Then, the line
@@
lies in the balanced manifold for all $t\in[0,1]$:
@@
    if $R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$.
  2. (2)
    if and only if
    (2.7) $(R_{p+1}-R_{p-1})M^{T}$
    are skew symmetric, where
    (2.8) $M=\tilde{\Lambda}^{-1}R_{p}\Lambda$
    for all $1\leq p\leq N-1$.
@@
Examples of when $A$ and $B$ satisfy this right-left rotation condition is if they are both symmetric or if they have the same singular vectors ($Q=\text{Id}$).
@@
This simplification can be shown by applying the theorem and expanding the geodesic formulas using that diagonal matrices commute. The limiting behavior as $N\to\infty$ (recognizing $X(t)$ is the expectation of $A^{S}B^{1-S}$ where $S$ is the average of NN Bernoulli(1t)(1-t) random variables) is
@@
We discuss this striking simplification more in Appendix A.
The two important tools we will use to prove Theorem 2.3 in Section 3.4 are (1) the characterization of straight lines on $\mathcal{M}$ (Lemma 2.2) and (2) an orthonormal basis for the horizontal space in $\mathcal{M}$ constructed by [10, 11, Theorem 34] (Lemma 3.10 and see [9] for an exposition). Straight lines are automatically geodesics on $(\mathcal{M},\iota)$ because $\iota$ is induced from the Euclidean metric on $\mathbb{M}_{d}^{N}$. We will then send a subset of these straight lines to new geodesics in $(\mathfrak{M}_{d},g^{N})$ to uncover the explicit formulas for geodesics as stated.
@@ 3.1. Geodesics Review
A geodesic between $A,B\in\mathcal{M}$ on $(\mathcal{M},g)$ is an extrema of the action
(3.1) $\mathcal{A}[x]=\int_{0}^{1}g_{x(t)}(\dot{x}(t),\dot{x}(t))dt$
@@
(3.2) $\mathcal{S}=\{x:[0,1]\to\mathcal{M}|x(0)=A,x(1)=B,x\in\mathcal{C}^{1}([0,1])\}.$
The minimizing $x$ is known to satisfy the Euler-Lagrange equations
@@
The dual of the Lagrangian optimization problem is the Hamiltonian formulation that reveals an equivalent set of equations of motion. Given a Lagrangian $L(x,\dot{x})$, the Hamiltonian is the Legendre transform of $L(x,\cdot)$
(3.4) $H(x,p)$ $=\sup_{\dot{x}}(\langle p,\dot{x}\rangle-L(x,\dot{x})).$
It is not hard to see that the change of variables is $p=\frac{\partial L}{\partial\dot{x}}$ when $\dot{x}$ is maximizing. The corresponding canonical equations to the Euler-Lagrange equations are
@@ 3.2. The Bures-Wasserstein Geometry
The Bures-Wasserstein geometry [5] is a special case of the DLN model when $N=2$ and $W_{1}=W_{2}^{T}$. The parameter space becomes $\mathfrak{M}_{d}$ whereas the observable space is $\mathbb{P}_{d}$. The map $\pi:\mathfrak{M}_{d}\to\mathbb{P}_{d}$ is a Riemannian submersion that maps
@@
It is also clear that $\mathcal{L}_{X}(Y)=XY+YX$. An important quantity in the Bures-Wasserstein geometry is the geometric mean of two matrices.
@@
As aforementioned, the Bures-Wasserstein geometry acts as a motivating example for our results. Specifically, we present a derivation of the explicit formulas for geodesics by solving the Hamiltonian equations of motion directly that we will generalize to the DLN (although we will not solve the equations explicitly).
@@
To begin, recall that $T_{X}\mathbb{P}_{d}\cong\text{Symm}_{d}$ and denote $Y=\dot{X}$. In order to construct the change of variable from Lagrangian into Hamiltonian coordinates, we must first compute $D_{Y}[L(X,Y)](H)$ where $H\in T_{X}\mathbb{P}_{d}$.
$D_{Y}\left[\frac{1}{2}\operatorname{Tr}(Y^{T}\mathcal{L}^{-1}_{X}(Y))\right](H)$ $=\frac{1}{2}\operatorname{Tr}(D_{Y}[Y^{T}\mathcal{L}^{-1}_{X}(Y)](H))$
$=\frac{1}{2}\operatorname{Tr}(H^{T}\mathcal{L}^{-1}_{X}(Y)+Y^{T}\mathcal{L}^{-1}_{X}(H))$
$=\frac{1}{2}(g_{X}(H,Y)+g_{X}(Y,H))$
(3.13) $=\operatorname{Tr}(\mathcal{L}^{-1}_{X}(Y)H)$
where the second to last equality holds because $H$ and $Y$ are in $T_{X}\mathbb{P}_{d}$ and the last equality follows from the cyclic property of trace. Thus, we obtain that $P=\mathcal{L}^{-1}_{X}(Y)$ and that $Y=\mathcal{L}_{X}(P)$. Now, we can compute the Hamiltonian using the cyclic property of trace and the definition of $P$ and $\mathcal{L}_{X}$ as
$H(X,P)$ $=\operatorname{Tr}(P^{T}\mathcal{L}_{X}(P))-\frac{1}{2}\operatorname{Tr}(\mathcal{L}_{X}(P)\mathcal{L}^{-1}_{X}(\mathcal{L}_{X}(P)))$
(3.14) $=\frac{1}{2}\operatorname{Tr}(P\mathcal{L}_{X}(P))=\operatorname{Tr}(P^{2}X).$
@@ 3.3. General Equations from Hamiltonian
We will now generalize the preceding calculation to the DLN. First, recall functional calculus on symmetric matrices [4, Chapter 2]. If $S=Q\Lambda Q^{T}$ is a symmetric matrix and $f:\mathbb{R}\to\mathbb{R}$ is a 1D functional, $f$ can be extended to the symmetric matrices as (abusing notation)
@@
$f^{[1]}_{\alpha}$ (Equation 2.1) now clearly comes from Equation 3.32 if $f(x)=x^{\alpha}$. We also have the following derivative computation.
@@
Rewriting $Z$ as $\tilde{Z}$ in the singular vector coordinates via the transformation $Z=U\tilde{Z}V^{T}$ nets the final simplification that gives the form in the Lemma statement. ∎
@@
(3.37) $\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\operatorname{Tr}([K\circ(A\Sigma)]M)=\operatorname{Tr}(\Sigma[K\circ M]A),$
(3.38) $\operatorname{Tr}([K\circ(\Sigma A)]M)$ $=\operatorname{Tr}([K\circ(A^{T}\Sigma)]M)=\operatorname{Tr}([K\circ M]\Sigma A).$
@@
Again, we will just prove for the first equation and note the second follows from almost identical steps. Let’s first prove that the first expression is equal to the last. We will approach by expanding into coordinates, following Einstein notation convention where repeated indices are summed.
$\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=K_{ij}\sigma_{i}(A^{T})_{ij}M_{ji}$
$=\sigma_{i}K_{ij}M_{ji}A_{ji}$
$=\sigma_{i}K_{ij}M_{ij}A_{ji}$
@@
$\operatorname{Tr}([K\circ(A\Sigma)]M)$ $=K_{ij}A_{ij}\sigma_{j}M_{ji}$
$=\sigma_{j}K_{ij}M_{ji}A_{ij}$
$=\sigma_{j}K_{ji}M_{ji}A_{ij}$
@@
Following similar steps to the Bures Wasserstein case, we can find the Hamiltonian on the manifold $(\mathfrak{M}_{d},g^{N})$ to be
@@
This first differential equation follows from exactly the same computation as the Bures-Wasserstein case (Equation 3.2) but with $\mathcal{L}^{-1}$ replaced by $\mathcal{A}^{-1}_{N,X}$. The only nontrivial equation is the second equation and the associated computation of $D_{X}[H(X,P)](Z)$. We begin by applying Lemma 3.5.
(3.42) $D_{X}[H(X,P)](Z)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}\left(T_{U}+T_{V}\right),$
@@
To clean up further, we also write $P$ in terms of the singular value coordinates by setting $\tilde{P}=U^{T}PV$, leaving just
(3.45) $T_{U}$ $=\tilde{P}^{T}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]\tilde{P}\Sigma^{\frac{2(p-1)}{N}}$
(3.46) $T_{V}$ $=\tilde{P}^{T}\Sigma^{\frac{2(N-p)}{N}}\tilde{P}\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right].$
Applying the trace identities from Lemma 3.6 to Equations 3.45 and 3.46, where we note $f^{[1]}_{\alpha}(\Sigma^{2})$ is symmetric, and substituting in $M_{1}$ and $M_{2}$ from Equation 2.3 reveal
@@
(3.49) $=\operatorname{Tr}\left(V\sum_{p=1}^{N}\left(\Sigma\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]+\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\Sigma\right)U^{T}Z\right)$

### v2

@@ Abstract.
We derive a general system of ODEs and associated explicit solutions in a special case for geodesics between full rank matrices in the deep linear network geometry. In the process, we find horizontal straight lines in the invariant balanced manifold that remain geodesics under Riemannian submersion.
@@ 1. Introduction
The deep linear network (DLN) is a phenomenological model of the training dynamics of neural networks [1, 2, 6, 11]. Given a depth $N\in\mathbb{Z}^{+}$ and width $d\in\mathbb{Z}^{+}$, the DLN in its simplest form consists of the “upstairs” (high-dimensional) parameter space of NN $d\times d$ matrices $\mathbb{M}_{d}^{N}$ and a “downstairs” (low-dimensional) observable space $\mathbb{M}_{d}$, linked by the product map $\phi:\mathbb{M}_{d}^{N}\to\mathbb{M}_{d}$ that sends
@@ 1.2. Balancedness and Riemannian Geometry
Interestingly, the DLN gradient flow (Equation 1.2) has invariant manifolds with associated metrics that illustrate the underlying geometric structure. One such invariant manifold is defined by the balancedness constraints
@@
for all $1\leq p\leq N-1$ [2, Theorem 1]. When all coordinates are also rank $d$, the manifold is called the balanced manifold and is denoted as $\mathcal{M}$. Note that when $\mathbf{W}\in\mathcal{M}$, $X=\phi(\mathbf{W})\in\mathfrak{M}_{d}$.
In $\mathcal{M}$, $W_{p}$ have the same singular values for all $1\leq p\leq N$ and the right singular vectors of $W_{p+1}$ align with the left singular vectors of $W_{p}$ for $1\leq p\leq N-1$. Thus, given the SVD of a matrix $X=Q_{N}\Sigma Q_{0}^{T}$, we can parameterize any $\mathbf{W}\in\mathcal{M}$ such that $\phi(\mathbf{W})=X$ using just
(1.5) $(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{0})\in\mathbb{R}_{+}^{d}\times O_{d}^{N+1},$
where $\Lambda=\Sigma^{\frac{1}{N}}$ and $W_{p}=Q_{p}\Lambda Q_{p-1}^{T}$ are obtained by the SVD for all $1\leq p\leq N$. Therefore, there exists a bijective map
@@
$g^{N}$ is crucial in the DLN because it can be used to express the training dynamics of the end-to-end matrix $X$ as a Riemannian gradient flow [3, Section 4]. Furthermore, these metrics in tandem provide a geometric perspective on the DLN model from [10, Theorem 8] and [13, Theorem 36], which we elaborate on in Section 3.4.1:
@@
Explicit formulas for geodesics in the Bures-Wasserstein geometry can be derived by identifying a Riemannian submersion structure from invertible matrices to the positive definite matrices [4, Section 4]. We present an alternate derivation of the geodesic formula that generalizes to the DLN, yielding the general ODEs for geodesics in $(\mathfrak{M}_{d},g^{N})$. Furthermore, Theorem 1.1 implies that the DLN also has a Riemannian submersion structure with $(\mathcal{M},\iota)$, $(\mathfrak{M}_{d},g^{N})$, and $\phi$, which we leverage to solve the general equations in a specific case.
@@ 2. Main Results
Our main results are a statement of ODEs for general geodesics in $(\mathfrak{M}_{d},g^{N})$ and explicit solutions when the singular vectors of the endpoints are related by a single matrix $Q\in O_{d}$. The solutions are obtained by finding a sufficient condition for horizontal straight lines in $\mathcal{M}$ in a manner similar to the classical theory of confocal quadrics [8].
@@
We discuss the origin of this matrix in Section 3.3. Let $\circ$ denote the Hadamard product between two matrices.
@@
$X(t)=U(t)\Sigma(t)V(t)^{T}$ is a geodesic between $A$ and $B$ on $(\mathfrak{M}_{d},g^{N})$ if and only if it satisfies the Dirichlet boundary value problem
@@
We discover these equations via direct computation in Hamiltonian coordinates (Section 3.3). Using Riemannian submersion, we are then able to derive explicit solutions to these equations for particular endpoints. As an important intermediate result, we find a sufficient condition for straight lines in $\mathcal{M}$.
Let $\mathbf{A}=\xi(\Lambda,Q_{N},\ldots,Q_{0})$ and $\mathbf{B}=\xi(\tilde{\Lambda},\tilde{Q}_{N},\ldots,\tilde{Q}_{0})$ be two endpoints in $\mathcal{M}$. Then, define
@@
Lemma 2.2 (Straight Lines on $\mathcal{M}$).
  1. (1)
    The line
@@
    is balanced for all $t\in[0,1]$ if and only if
    (2.7) $(R_{p+1}-R_{p-1})\left(\tilde{\Lambda}^{-1}R_{p}\Lambda\right)^{T},$
    is skew symmetric for all $1\leq p\leq N-1$.
@@
    If $R_{p}=R_{p-1}$ for all $1\leq p\leq N$, $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in[0,1]$.
@@
Examples of when $A$ and $B$ satisfy this right-left rotation condition are if they are both symmetric or if they have the same singular vectors ($Q=\text{Id}$).
@@
This simplification can be shown by applying the theorem and expanding the geodesic formulas using the commutativity of diagonal matrices. The limiting behavior as $N\to\infty$ (recognizing $X(t)$ as the expectation of $A^{S}B^{1-S}$ where $S$ is the average of NN Bernoulli(1t)(1-t) random variables) is
@@
The two important tools we will use to prove Theorem 2.3 in Section 3.4 are (1) the sufficient condition for straight lines on $\mathcal{M}$ (Lemma 2.2) and (2) an orthonormal basis for the horizontal space in $\mathcal{M}$ (Lemma 3.11 and [10, 13]). Straight lines are automatically geodesics on $(\mathcal{M},\iota)$ because $\iota$ is induced from the Euclidean metric on $\mathbb{M}_{d}^{N}$. We will then send a subset of these straight lines to new geodesics in $(\mathfrak{M}_{d},g^{N})$ to uncover the explicit formulas for geodesics as stated.
@@ 3.1. Geodesics Review
For more details on geodesics, we refer the reader to [7]. Recall the following standard definitions and results.
@@
A geodesic between $A,B\in\mathcal{X}$ on the Riemannian manifold $(\mathcal{X},g)$ is an extrema of the action
(3.1) $\mathcal{A}[x]=\int_{0}^{1}L(x,\dot{x})dt,\quad L(x,\dot{x})=g_{x(t)}(\dot{x}(t),\dot{x}(t))$
@@
(3.2) $\mathcal{S}=\{x:[0,1]\to\mathcal{X}|x(0)=A,x(1)=B,x\in\mathcal{C}^{1}([0,1])\}.$
$L$ is referred to as the Lagrangian.
Theorem 3.2 (Euler-Lagrange and Hamiltonian Equations).
A geodesic $x$ satisfies the Euler-Lagrange equations:
@@
and the dual Hamiltonian equations
@@
where $H$ is the Legendre transform of $L(x,\dot{x})$
(3.6) $H(x,p)=\sup_{\dot{x}}(\langle p,\dot{x}\rangle-L(x,\dot{x})).$
with change of variables $p=\frac{\partial L}{\partial\dot{x}}$.
For a Riemannian manifold $(\mathcal{X},g)$, the Hamiltonian $H$ associated with Lagrangian $L(x,\dot{x})=\frac{1}{2}g(\dot{x},\dot{x})$ is always $H(x,p)=\frac{1}{2}g^{-1}(p,p)$, where $g^{-1}$ is the inverse metric tensor acting on cotangent vectors.
@@ 3.2. The Bures-Wasserstein Geometry
The Bures-Wasserstein geometry [4] is a special case of the DLN model when $N=2$ and $W_{1}=W_{2}^{T}$ that will serve as a motivating example for the derivation of Theorem 2.1. The parameter space is $\mathfrak{M}_{d}$ whereas the observable space is $\mathbb{P}_{d}$. The map $\pi:\mathfrak{M}_{d}\to\mathbb{P}_{d}$ is a Riemannian submersion that maps
@@
The inverse metric tensor is $\mathcal{L}_{X}(Y)=XY+YX$. An important quantity in the Bures-Wasserstein geometry is the geometric mean of two matrices.
@@
Our main computation in the Bures-Wasserstein geometry is a derivation of the explicit formulas for geodesics by solving the Hamiltonian equations of motion directly.
@@
To begin, recall that $T_{X}\mathbb{P}_{d}\cong\text{Symm}_{d}$ and denote $Y=\dot{X}$. We consider the Hamiltonian system, which from the inverse metric tensor $\mathcal{L}$ is
(3.13) $H(X,P)$ $=\frac{1}{2}\operatorname{Tr}(P\mathcal{L}_{X}(P))=\operatorname{Tr}(P^{2}X).$
@@
The uniqueness of this solution also follows from Picard’s theorem, as the right hand sides of Equation 3.11 are smooth. ∎
@@ 3.3. General Equations from Hamiltonian
We will now generalize the preceding calculation to the DLN. First, recall functional calculus on symmetric matrices [5, Chapter 2]. If $S=Q\Lambda Q^{T}$ is a symmetric matrix and $f:\mathbb{R}\to\mathbb{R}$ is a 1D functional, $f$ can be extended to the symmetric matrices as (abusing notation)
@@
Equation 2.1 for $f^{[1]}_{\alpha}$ is the special case of Equation 3.31 when $f(x)=x^{\alpha}$. We also have the following derivative computation.
@@
The final simplification comes from expanding using $Z=U\tilde{Z}V^{T}$. ∎
@@
(3.36) $\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\operatorname{Tr}([K\circ(A\Sigma)]M)=\operatorname{Tr}(\Sigma[K\circ M]A)$
@@
First, we prove that the first expression is equal to the last. We will approach by expanding into coordinates.
$\operatorname{Tr}([K\circ(\Sigma A^{T})]M)$ $=\sum_{i}\sum_{j}K_{ij}\sigma_{i}(A^{T})_{ij}M_{ji}$
$=\sum_{i}\sum_{j}\sigma_{i}K_{ij}M_{ji}A_{ji}$
$=\sum_{i}\sum_{j}\sigma_{i}K_{ij}M_{ij}A_{ji}$
@@
$\operatorname{Tr}([K\circ(A\Sigma)]M)$ $=\sum_{i}\sum_{j}K_{ij}A_{ij}\sigma_{j}M_{ji}$
$=\sum_{i}\sum_{j}\sigma_{j}K_{ij}M_{ji}A_{ij}$
$=\sum_{i}\sum_{j}\sigma_{j}K_{ji}M_{ji}A_{ij}$
@@
The Hamiltonian on the manifold $(\mathfrak{M}_{d},g^{N})$ is again given by the inverse metric tensor:
@@
The first equation follows directly from the Hamiltonian. The second equation requires computation of $D_{X}[H(X,P)](Z)$. We begin by applying Lemma 3.6.
(3.40) $D_{X}[H(X,P)](Z)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}\left(T_{U}+T_{V}\right)=\frac{1}{2}\sum_{p=1}^{N}\operatorname{Tr}(T_{U})+\operatorname{Tr}(T_{V}),$
@@
To clean up further, we can write $P$ in terms of the singular value coordinates by setting $\tilde{P}=U^{T}PV$ and using the cyclic property of trace to see that
(3.43) $\operatorname{Tr}(T_{U})$ $=\operatorname{Tr}\left(\tilde{P}^{T}\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}^{T}+\tilde{Z}\Sigma\right)\right]\tilde{P}\Sigma^{\frac{2(p-1)}{N}}\right)$
(3.44) $\operatorname{Tr}(T_{V})$ $=\operatorname{Tr}\left(\tilde{P}^{T}\Sigma^{\frac{2(N-p)}{N}}\tilde{P}\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ\left(\Sigma\tilde{Z}+\tilde{Z}^{T}\Sigma\right)\right]\right).$
Applying the identities from Lemma 3.7 to Equations 3.43 and 3.44 ($f^{[1]}_{\alpha}(\Sigma^{2})$ is symmetric) and substituting in $M_{1}$ and $M_{2}$ from Equation 2.3 shows
@@
(3.47) $=\operatorname{Tr}\left(V\sum_{p=1}^{N}\left(\Sigma\left[f_{\frac{N-p}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{1}\right]+\left[f_{\frac{p-1}{N}}^{[1]}\left(\Sigma^{2}\right)\circ M_{2}\right]\Sigma\right)U^{T}Z\right).$